\begin{thebibliography}{AAAA}
%
%\bibitem[Ch]{Cheah}
%J.~Cheah, {\em The Hogde polynomial of the Fulton--MacPherson compactification of configuration spaces},
% Amer.\ J.\
%Math.\ {\bf 118} (1996), 963--977.
%

%\bibitem[A]{Almgren} F. Almgren, {\sl The homotopy
%groups of the integral cycle groups.} Topology 1 1962 257--299.

\bibitem[B]{Bloch} S. Bloch,
{\sl Algebraic cycles and higher $K$-theory.}
Adv. in Math. 61 (1986), no. 3, 267--304.


\bibitem[C]{Cheeger} J. Cheeger, {\sl Multiplication of differential characters.} Symposia
Mathematica, Vol. XI (Convegno di Geometria, INDAM, Rome, 1972), pp.
441--445. Academic Press, London, 1973.

\bibitem[CS]{Cheeger-Simons}
J. Cheeger; J. Simons {\sl Differential characters and geometric
invariants.} Geometry and topology (College Park, Md., 1983/84),
50--80, Lecture Notes in Math., 1167, Springer, Berlin, 1985.


\bibitem[EV]{Esnault-Viehweg} H. Esnault; E. Viehweg,
{\sl Deligne-Be\u\i linson cohomology.} Be\u\i linson's conjectures
on special values of $L$-functions, 43--91, Perspect. Math., 4,
Academic Press, Boston, MA, 1988.


\bibitem[F]{Friedlander1} E. Friedlander,  {\sl Algebraic cycles, Chow
varieties, and Lawson homology.}  Compositio Math. 77 (1991), no. 1,
55--93.

\bibitem[FG]{Friedlander-Gabber} E. Friedlander and O. Gabber, {\sl Cycle
spaces and intersection theory. Topological methods in modern
mathematics} (Stony Brook, NY, 1991), 325--370, Publish or Perish,
Houston, TX, 1993.


\bibitem[FL]{Friedlander-Lawson}E. Friedlander and  B. Lawson,
{\sl A theory of algebraic cocycles.}
Ann. of Math. (2) 136 (1992), no. 2, 361--428.

\bibitem[FL2]{Friedlander-Lawson2}E. Friedlander and  B. Lawson,
{\sl Duality relating spaces of algebraic cocycles and cycles.} (English summary)
Topology 36 (1997), no. 2, 533--565.


\bibitem[FM]{Friedlander-Mazur} E. Friedlander and B. Mazur,
{\sl Filtrations on the homology of algebraic varieties. With an
appendix by Daniel Quillen.}  Mem. Amer. Math. Soc. 110 (1994), no.
529, x+110 pp.


\bibitem[GS]{Gillet-Soule} H. Gillet and C. Soul\'{e} {\sl Arithmetic Chow
groups and differential characters.} Algebraic $K$-theory:
connections with geometry and topology (Lake Louise, AB, 1987),
29--68, NATO Adv. Sci. Inst. Ser. C Math. Phys. Sci., 279, Kluwer
Acad. Publ., Dordrecht, 1989.

\bibitem[G]{Griffiths} P. Griffiths, {\sl On the periods of certain
rational integrals I, II}. Ann. of Math. (2) 90(1969), 460-495;
ibid. (2) 90(1969) 496--541.

\bibitem[Ha]{Harris} B. Harris,  {\sl Differential characters and the
Abel-Jacobi map.} Algebraic $K$-theory: connections with geometry
and topology (Lake Louise, AB, 1987), 69--86, NATO Adv. Sci. Inst.
Ser. C Math. Phys. Sci., 279, Kluwer Acad. Publ., Dordrecht, 1989.

\bibitem[HL1]{Harvey-Lawson1} R. Harvey and B. Lawson {\sl From Sparks to
Grundles - Differential Characters.} (English summary)
Comm. Anal. Geom. 14 (2006), no. 1, 25--58.


\bibitem[HL2]{Harvey-Lawson2}  R. Harvey and  B. Lawson,  {\sl D-bar
sparks}.    (English summary)
Proc. Lond. Math. Soc. (3) 97 (2008), no. 1, 1--30.

\bibitem[HLZ]{Harvey-Lawson-Zweck} R. Harvey; B. Lawson and J. Zweck, {\sl
The de Rham-Federer theory of differential characters and character
duality.}  Amer. J. Math. 125 (2003), no. 4, 791--847.


\bibitem[HP]{Harvey-Polking} R. Harvey; J. Polking {\sl Fundamental solutions in complex
analysis.} I and II. The Cauchy-Riemann operator. Duke Math. J. 46
(1979), no. 2, 253--340.

\bibitem[H]{aj} W. Hu, {\sl  Generalized Abel-Jacobi map on Lawson
Homology.} arXiv: math.AG/0608294. To appear in  Amer. J. Math.



\bibitem[L1]{Lawson1} B. Lawson, {\sl Algebraic cycles and homotopy
theory.} Ann. of Math. {\bf 129}(1989), 253-291.

\bibitem[L2]{Lawson2} B. Lawson,  {\sl Spaces of algebraic cycles.} pp.
137-213 in Surveys in Differential Geometry, 1995 vol.2,
International Press, 1995.


\bibitem[LF]{Lima-Filho3}P. Lima-Filho,
{\sl Completions and fibrations for topological monoids.}
Trans. Amer. Math. Soc. 340 (1993), no. 1, 127--147.


\bibitem[KLM]{KLM}
M. Kerr,  J. Lewis and  S. M\"{u}ller-Stach,
{\sl The Abel-Jacobi map for higher Chow groups.}  (English summary)
Compos. Math. 142 (2006), no. 2, 374--396.

\bibitem[Pe]{Peters} C. Peters, {\sl  Lawson homology for varieties with
small Chow groups and the induced filtration on the Griffiths
groups.}  Math. Z. 234 (2000), no. 2, 209--223.



\bibitem[W]{Walker}
M. Walker
{\sl The morphic Abel-Jacobi map.} (English summary)
Compos. Math. 143 (2007), no. 4, 909--944.
\end{thebibliography}

\documentclass[10pt]{amsart}
\usepackage{amssymb,latexsym,amsmath}
\usepackage{graphics}                 % Packages to allow inclusion of graphics
\usepackage{color}                    % For creating coloured text and background
\usepackage{hyperref}                 % For creating hyperlinks in cross references
\usepackage[all]{xy}

\newtheorem{definition}{Definition}[section]
\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{remark}[theorem]{Remark}
\newtheorem{example}[theorem]{Example}

\def\C{{\mathbb{C}}}

\def\R{{\rm R}}
\def\rS{{\rm S}}
\def\Z{{\mathbb{Z}}}
\def\Ch{{\rm Ch}}

\def\H{{\bf H}}

\def\D{{\mathcal{D}}}
\def\mC{{\mathcal{C}}}
\def\E{{\mathcal{E}}}
\def\cR{{\mathcal{R}}}
\def\cZ{{\mathcal{Z}}}

\title{A map from Lawson homology to Deligne Cohomology}
\author{Wenchuan Hu}
\date{September 29, 2009}
\address{Institute for Advanced Study\\
Einstein Drive\\
Princeton, N.J. 08540\\
U.S.A. }

\email{wenchuan@math.ias.edu}

\begin{document}

\begin{abstract}A natural  Abel-Jacobi type map from
Lawson homology to Deligne cohomology  for smooth
complex projective varieties is constructed by using the Harvey-Lawson ``spark"
complexes. We also compare this to  Abel-Jacobi type constructions by others.
\end{abstract}

\maketitle
%\pagestyle{myheadings}
% \markright{}

\tableofcontents

\section{Introduction} Let $X$ be a complex projective variety of
dimension $n$.  Let ${\mC}_p(X)$ be the space of all effective algebraic $p$-cycles on $X$,
 i.e., an element $c\in {\mC}_p(X)$  is a finite sum $c=\sum n_i V_i$, where $n_i$ is a
nonnegative integer and $V_i$ is an irreducible variety of dimension $p$.
Let ${\cZ}_p(X)$ be the space of all algebraic $p$-cycles on $X$.

The \textbf{Lawson homology} $L_pH_k(X)$ of
$p$-cycles is defined by
$$L_pH_k(X) := \pi_{k-2p}({\cZ}_p(X))~ \hbox{for $k\geq 2p\geq 0$},$$
where ${\cZ}_p(X)$ is provided with a natural topology so that
it is an abelian topological group (cf. \cite{Friedlander1},
\cite{Lawson1} and \cite{Lawson2}). For general background, the reader is referred to
\cite{Lawson2}.

Friedlander and Mazur \cite{Friedlander-Mazur} showed that there
are  natural maps, called \textbf{cycle class maps} $\Phi_{p,k}:L_pH_{k}(X)\rightarrow H_{k}(X)$
from Lawson homology to singular homology. Denote by $L_pH_{k}(X)_{hom}$ the kernel of $\Phi_{p,k}$,
i.e., $L_pH_{k}(X)_{hom}={\rm ker}\{\Phi_{p,k}:L_pH_{k}(X)\rightarrow H_{k}(X)\}.$

Assume that $X$ is a smooth complex manifold. Let $\Omega_X^k$
the sheaf of holomorphic $k$-form on $X$. The \textbf{Deligne
complex of level p} is the complex of sheaves
$$\underline{\Z}_{\D}(p):0\rightarrow\Z(p)\rightarrow\Omega_X^0\rightarrow
\Omega_X^1\rightarrow\Omega_X^2\rightarrow\cdots\rightarrow\Omega_X^{p-1}\rightarrow
0,
$$
where $\Z(p):=(2i\pi)^p\cdot\Z\subset \C$.

The \textbf{Deligne cohomology} of $X$ in level $p$ is defined as the
hypercohomology of this complex
$ H^*_{\D}(X,\Z(p)):= \H^*(X,\underline{\Z}_{\D}(p)).$

For more details on Deligne cohomology, the reader is referred to
 \cite{Esnault-Viehweg}.

We  define a natural map from
Lawson homology to the corresponding Deligne cohomology
using the theory of differential characters  introduced by
Cheeger in \cite{Cheeger} and  the theory of D-bar sparks developed
by Harvey and Lawson in \cite{Harvey-Lawson1} and
\cite{Harvey-Lawson2}.
 This theory of differential characters has been developed in \cite{Cheeger-Simons}, \cite{Gillet-Soule}, \cite{Harris}
and systematically generalized by Harvey, Lawson and Zweck in
\cite{Harvey-Lawson-Zweck}.
This map is a generalization
of the Abel-Jacobi map for Lawson homology defined by the author in \cite{aj}.

 The main result in this paper is the following

\begin{theorem}
\label{sec:T7Th1.1} Let $X$ be a smooth complex projective variety
of dimension $n$. We have a well-defined homomorphism
$$\hat{a}_X: L_pH_{k+2p}(X)\rightarrow
H^{2(n-p)-k}_{\D}(X,\Z(n-p-k-1)),$$ given by
$$\hat{a}_X([f])=\widehat{a_f}
$$
which coincides with the generalized Abel-Jacobi map defined in
\cite{aj} when $\hat{a}$ is restricted on $L_pH_k(X)_{hom}$ and the
projection of the image of $\hat{a}$ under $\delta_2$ is the natural
map $\Phi_{p,k}$.
\end{theorem}

If there is no confusion, $\hat{a}_X$ is simply denoted by $\hat{a}$.
The notations $\widehat{a_f}$ and $\delta_2$ will be defined below.

We also obtain functorial properties for $\hat{a}$, that is, it commutes with
projective morphisms between smooth complex projective varieties (cf. Proposition \ref{prop3.10}). Moreover, our construction is also applied to higher Chow groups and we obtain a natural homomorphism from higher Chow groups to Deligne cohomology (cf. Proposition \ref{prop4.4}). Our maps on higher Chow groups factors those on Lawson homology (cf. Corollary \ref{cor4.5}). However, they seem to be different from those (called regulator maps) given by Bloch \cite{Bloch}.

\section{Sparks and differential characters}\label{sec2}
In this section we review the necessary  background on  materials of sparks and
differential characters for our construction in section \ref{sec3}. For details, see
\cite{Harvey-Lawson-Zweck}, \cite{Harvey-Lawson1} and
\cite{Harvey-Lawson2}. In this section $X$ denotes a smooth manifold unless otherwise noted.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\begin{definition}
Set ${\E}^k(X):=$ the space of smooth differential forms
$k$-form on $X$ with $C^{\infty}$-topology; ${\D}^k(X):=$\{$\phi\in {\E}^k(X) |$ supp($\phi$) is compact\}.
We say the space of \emph{currents of degree $k$ (and dimension
$n-k$)} on $X$, it means the topological dual space
${\D'}^k(X)\equiv {\D'}_{n-k}(X):=\{{\D}^{n-k}(X)\}'$.
Set

${\cR}^k(X):=$ the locally {\rm rectifiable} currents of degree
k(dimension $n-k$) on $X$;

${\mathcal IF}^k(X):=$ the locally  {\rm integrally flat} currents of
degree k on $X$;
and

 ${\mathcal I}^k(X):=$ the locally  {\rm integral} currents of degree $k$ on $X$.
\end{definition}

 The following notation was firstly given in
\cite{Harvey-Lawson-Zweck}:

\begin{definition}
The space of \emph{{sparks}} of degree $k$ on $X$ is defined to
be
\footnotesize{
$$ {\mathcal{S}}^k(X):=\{s\in {\D'}^k(X)| da =\phi-R, \hbox{
where $\phi\in {\E}^{k+1}(X)$ and $R\in
{\mathcal IF}^{k+1}(X)$}\}.
$$
}
\end{definition}

\begin{definition}
For each integer $k$, $0\leq k\leq n$, the de
Rham-Federer characters of degree $k$  id defined to  be the quotient
$$ \widehat{\H}^k(X):={\mathcal{S}}^k(X)/\{d{\D'}^{k-1}(X)\}+{\mathcal IF}^{k}(X)\}
$$
\end{definition}

The equivalence class in $\widehat{\H}^k(X)$ of a spark $a\in
{\mathcal{S}}^k(X)$ will be denoted by $\hat{a}$.

It has been proved that $\phi$ and $R$ in the decomposition of $da$
above is unique (cf. \cite{Harvey-Lawson-Zweck}, Lemma 1.3). Moreover,
there are two well-defined surjective maps:
$$\delta_1:\widehat{\H}^k(X)\rightarrow {Z}_0^{k+1}(X);\quad
\delta_1(\hat{a})=\phi
$$
and
$$\delta_2:\widehat{\H}^k(X)\rightarrow
H^{k+1}(X,\Z); \quad \delta_2(\hat{a})=[R],
$$
where ${Z}_0^{k+1}(X)$ denotes the  lattice of smooth
$d$-closed, degree $k+1$ forms on $X$ with integral periods.

Now we can give the definition of Riemannian Abel-Jacobi map. Let
$X$ be compact Riemannian manifold. Any current $R$ on $X$, has  a
Hodge decomposition (cf. \cite{Harvey-Polking})
$$ R=H(R)+dd^*G(R)+d^*dG(R)
$$
where $H$ is harmonic projection and $G$ is the Green operator.
Also recall that $d$ commutes with $G$, so that if $R$ is a cycle,
then $dG(R)=0$. For $R\in {\mathcal IF}^{k+1}(X)$, set
$a(R):=-d^*G(R)$ then $$da(R)=H(R)-R$$ i.e., $a(R)$ is a {\sl
Hodge spark}. Let $\widehat{a(R)}\in \widehat{\H}^k(X)$ denote the
differential character corresponding to the Hodge spark $a(R)$.
Set
$${\rm Jac}^k(X):=H^k(X;\mathbb{R})/H^k_{\rm free}(X;\Z); \quad{\mathcal B}^{k+1}(X):=d{\mathcal IF}^k(X)
$$
then we have a well-defined map
$$\hat{a}:{\mathcal B}^{k+1}(X)\rightarrow {\rm Jac}^k(X)
$$
which is called the $k$-th {\sl Riemannian Abel-Jacobi map}.

In \cite{Harvey-Lawson1},  the concept of \textbf{homological spark
complex} and its associated \textbf{group of homological spark
classes} are given. In \cite{Harvey-Lawson2}, a generalized version
of homological spark complex has been defined.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\begin{definition}
A \textbf{homological spark complex} is a triple of cochain
complexes $(F^*,E^*,I^*)$ together with morphisms
$$\Psi:I^*\rightarrow F^*\supset E^*
$$
such that:
\begin{enumerate}
 \item $\Psi(I^k)\cap E^k=\{0\}$ for $k>0$,
 \item $H^*(E)\cong H^*(F)$, and
 \item $\Psi:I^0\rightarrow F^0$ is injective.
 \end{enumerate}
\end{definition}

\begin{definition}
\label{sec:T7def2.5} In a given spark complex $(F^*,E^*,I^*)$
a \textbf{spark of degree} k is a pair
 $$ (a,r)\in F^k\oplus I^{k+1}
$$
which satisfies the spark equation
\begin{enumerate}
 \item $da=e-\Psi(r)$ for some $e\in E^{k+1}$, and
 \item $dr=0$.
 \end{enumerate}
The group of sparks of degree k is denoted by ${\mathcal S}^k={\mathcal S}^k(F^*,E^*,I^*)$.
\end{definition}

\begin{definition}
Two sparks $(a,r),(a',r')\in {\mathcal S}^k(F^*,E^*,I^*)$ are
equivalent if there exists a pair
$$ (b,s)\in F^{k-1}\oplus I^k
$$
such that
\begin{enumerate}
 \item $a-a'=db+\Psi(s)$, and
 \item $r-r'=-ds$.
 \end{enumerate}
The set of equivalence classes is called the \textbf{group of spark
classes of degree} k associated to the given spark complex and will
be denoted by $\widehat{\H}^k(F^*,E^*,I^*)$ or simply
$\widehat{\H}^k(F)$.
\end{definition}

As usual, let $Z^k(E)=\{e\in E^k| de=0\}$ and set
$$Z^k_I(E):=\{e\in Z^k(E)|[e]=\Psi_*(\rho)~ \hbox{for some $\rho \in H^k(I)$}\}
$$
where $[e]$ denotes the class of $e$ in $H^k(E)$ (note that $de=0$).
The following lemma was proved in \cite{Harvey-Lawson-Zweck}:

\begin{lemma} \label{sec:T7L2.1}
There exist well-defined surjective homomorphisms:
$$
\delta_1: \widehat{\H}^k(F)\rightarrow Z^k_I(E)\quad and\quad
\delta_2: \widehat{\H}^k(F)\rightarrow H^{k+1}(I)
$$
given on any representing spark $(a, r)\in {\mathcal S}^k$ by
$$
\delta_1(a,r)=e \quad and \quad\delta_2(a,r)=[r]
$$
where $da=e-\Psi(r)$ as in Definition \ref{sec:T7def2.5}.
\end{lemma}

The following  example is the
main object which will be dealt with in the next section.
\begin{example}\label{eg2.2}
Now let $X$ be a smooth complex projective variety of  dimension $n$. Set
$$
F^m={\D}'^{m}(X,q):=\oplus_{r+s=m,r< q} {\D}'^{r,s}(X)
\quad {\rm and}\quad \bar{d}=\Psi\circ d
$$
where
$$
\Psi:{\D}'^{m}(X)\rightarrow{\D}'^{m}(X,q)
$$
 is the projection
$\Psi(a)=a^{0,m}+a^{1,m-1}+\cdots+a^{q-1,m-q+1}$.
$$
E^m={\E}^{m}(X,q):=\oplus_{r+s=m,r<q} {\E}^{r,s}(X)
\quad {\rm and}\quad \bar{d}=\Psi\circ d
$$
 and
$$
I^m={\mathcal I}^m(X)
$$

It has been shown in \cite{Harvey-Lawson2} that the above
triple $(F^*,E^*,I^*)$ is a homological spark complex . The group of
associated spark classes in degree $m$ will be denoted by
$\widehat{\H}^m(X,q)$. From this homological spark complex, one has
$$
\ker(\delta_1)=H^{m+1}_{\D}(X,\Z(q)).
$$

Denote by $H^*(X,q)$ the cohomology $H^*(E,q)\cong H^*(F,q)$ and set
$H^{*}_{\Z}(X,q):={\rm Image}\{\Psi_*:H^*(X,\Z)\to H^*(X,q)\}$.
\end{example}

%%%%%%%%%%%%%%%sect33333333333333333333333333333333333333333333333333333333

 \section{The construction of the map}\label{sec3}
 In this section, $X$ denotes a smooth projective variety.
Let ${\cZ}_p(X)$ be the space of algebraic $p$-cycles with a
natural topology  and a base point, i.e., the `null' $p$-cycle
(cf. \cite{Lawson1}, \cite{Lawson2} or \cite{Friedlander1}).
Let $\Omega^k{\cZ}_p(X)$ be the $k$-th iterated loop space with the given base
point. Explicitly, we write ${\rS}^k$ as $\R^k\cup {\infty}$ and we have
$$\Omega^k{\cZ}_p(X)=\{f:{\rS}^k\rightarrow {\cZ}_p(X)|\hbox{ $f$
 is continuous with $f(\infty)=0$}\}.$$

For $k=0$, an element $f\in \Omega^0{\cZ}_p(X)$ is the difference of two map $f_1,f_2$, where
$f_i:{\rS}^0\rightarrow {\mC}_p(X)$ are maps from ${\rS}^0$ to the space of effective $p$-cycles
${\mC}_p(X)$ for $i=1,2$  (cf. \cite{Lima-Filho3} or \cite{Friedlander-Gabber}). Therefore
$\Omega^0{\cZ}_p(X)$ can be identified with the space containing differences of effective $p$-cycles,
i.e., ${\cZ}_p(X)$. The map $f$ determines an integral current $c_f=f_1(s_0)-f_2(s_0)$, where $s_0\in {\rS}^0-\{\infty\}$.
In fact $c_f$ is a closed integral current since it is an algebraic cycle on $X$. By definition,
the homology class  of $c_f$ depends only on the homotopy class  $[f]\in \pi_0{\cZ}_p(X)\cong L_pH_{2p}(X)$ of $f$,
where the isomorphism was given in \cite{Friedlander1}.
Note that $c_f\in {\D}'_{p,p}(X)$.

For $k>0$ and a continuous map $f:{\rS}^k\rightarrow {\cZ}_p(X)$,  we
can find a map $g:{\rS}^k\rightarrow {\mC}_p(X)$ such that $g$ is
homotopic to $f$ (cf. \cite{Lima-Filho3} or \cite{Friedlander-Gabber}) and
 $g$ is piecewise linear with regard to a
triangulation of ${\mC}_p(X)$.
 Hence one can define an integral current
$c_g$ over $X$. Note that this current $c_g$ is actually a cycle, i.e., $d(c_g)=0$.
 Moreover, the homology class of $c_g$ depends only on the homotopy class of $g$. The detail of the
construction of $c_g$ from the map $g:{\rS}^k\rightarrow {\cZ}_p(X)$ can be found
in   section 3 of \cite{aj}.

In any case,  we obtain an integral cycle $c_f$ from an element $f\in \Omega^k{\cZ}_p(X)$. In the following
of this section we will construct an element in a suitable Deligne cohomology group from the integral cycle $c_f$.

Consider Example \ref{eg2.2}  with
$m=2(n-p)-k-1,q=n-p-k-1$. All the following argument will focus on
this homological spark complex.

\begin{definition}\label{def3.1}
 Set $a_f:= -\Psi (d^*G(c_f))$. Then $(a_f,c_f)$ is called the
Hodge spark of the map $f:{\rS}^k\rightarrow {\cZ}_p(X)$, where $G$ is the Green's operator.
Let $$\widehat{a_f}\in \widehat{\H}^{2(n-p)-k-1}(X,n-p-k-1)$$ be the
differential character corresponding to the Hodge spark
$(a_f,c_f)$.
\end{definition}

\begin{remark}
 In Definition \ref{def3.1}, we need to choose a Riemannian metric of $X$. However,
$\widehat{a_f}$ is independent of the choice of such a metric (cf. \cite{Harvey-Lawson-Zweck}, p. 830.).
\end{remark}

\begin{lemma}
{\label{sec:T7L3.1} If $f\in \Omega^k{\cZ}_p(X)$, then
$\delta_1(\widehat{a_f})=0\in
\widehat{\H}^{2(n-p)-k-1}(X,n-p-k-1)$.}
\end{lemma}

 \begin{proof} By definition of $\delta_1$, we have $\delta_1(\widehat{a_f})=\widehat{\Psi(H(c_f))}$,
where $H(c_f)$ is the harmonic part of $c_f$. Note that
$H(c_f)$ and $c_f$ are of the same $(*,*)$-type. Moreover, from
the construction of $c_f$, we have
 $$c_f\in
\bigoplus_{r+s=k,|r-s|\leq k} {\D}'_{p+r,p+s}(X)$$
(cf. \cite{aj}, Remark 3.3)

 By the type
reason, the projection of $H(c_f)$ under $\Psi$ on
$$\bigoplus_{r+s=k,r> k+1} {\D}'_{p+r,p+s}(X)
$$ is zero. This is exactly the image of $\widehat{a_f}$ under
$\delta_1$.
\end{proof}

Hence we get an element $\widehat{a_f}\in \ker(\delta_1)=
H^{2(n-p)-k}_{\D}(X,\Z(n-p-k-1))$. Therefore we get a map
$$\hat{a}: \Omega^k{\cZ}_p(X) \rightarrow
H^{2(n-p)-k}_{\D}(X, \Z(n-p-k-1))$$
which is defined by $\hat{a}(f)=\widehat{a_f}$.

\iffalse
\begin{remark}
{By the argument above, we have actually defined an element
$$\widehat{a_f}\in  H^{2(n-p)-k}_{\D}(X,\Z(n-p-k)).$$}
\end{remark}
\fi

If the map $f$ is homotopically trivial, then we have the following:
\begin{lemma}
 \label{sec:T7L3.2} If $f:{\rS}^k\rightarrow {\cZ}_p(X)$ is
homotopically trivial, then $\hat{a}(f)=0$.
\end{lemma}

\begin{proof}
Let $F:D^{k+1}\rightarrow {\cZ}_p(X)$ be an extension of
$f$, i.e., $F|_{\partial(D^{k+1})}=f$. Let $c_F$ be the current over
$X$ defined by $F$. As showed in Lemma \ref{sec:T7L3.1}, $c_F$ is an
integral current with boundary $\partial(c_F)=c_f$. By
Corollary 12.11 in \cite{Harvey-Lawson-Zweck}, we have
$\widehat{a_f}=\widehat{ H(\Psi(c_F))}$. Since
$$c_F\in\bigoplus_{r+s=k+1,|r-s|\leq k+1} {\D}'_{p+r,p+s}(X),$$
the projection of $H(c_F)$ under $\Psi$ on
$$\bigoplus_{r+s=k,r> k+1} {\D}'_{p+r,p+s}(X)$$
is zero. Note that $\Psi$ commutes with the Laplace operator, we
have $\widehat{a_f}=0$.
\end{proof}

By the Lemma \ref{sec:T7L3.1} and Lemma \ref{sec:T7L3.2}, we have a
well-defined map
$$\hat{a}: L_pH_{k+2p}(X)\rightarrow
H^{2(n-p)-k}_{\D}(X,\Z(n-p-k-1)),$$ given by
$$\hat{a}([f])=\widehat{a_f}.
$$
Recall that the Deligne cohomology can be written as the middle
part of a short exact sequence
{\tiny
$$0\rightarrow \frac{H^{2(n-p)-k-1}(X,n-p-k-1)}{H^{2(n-p)-k-1}(X,\Z)}\rightarrow
H^{2(n-p)-k}_{\D}(X,\Z(n-p-k-1))\rightarrow
\ker(\Psi_*)\rightarrow 0,
$$}
where
$$H^{2(n-p)-k-1}(X,n-p-k-1)=\bigg\{\bigoplus_{r+s=2(n-p)-k-1,r<n-p-k-1}H^{r,s}(X)\bigg\}$$
 and
$$\ker(\Psi_*)= H^{2(n-p)-k}(X,\Z)\cap \bigg\{\bigoplus_{r+s=2(n-p)-k,|r-s|\leq k+2}
H^{r,s}(X)\bigg\}.
$$

\begin{proposition}
\label{sec:T7Prop3.1} The restriction of the above map to
$$L_pH_{k+2p}(X)_{hom}:= \ker\{ \Phi_{p,k+2p}:L_pH_{k+2p}(X)\rightarrow H_{k+2p}(X,\Z)\}$$
is the generalized Abel-Jacobi map defined in \cite{aj} if we
identify the $H^{r,s}(X)$ with $\{H^{n-r,n-s}(X)\}^*$ for all $0\leq
r,s\leq n$ and $H^{q}(X,\Z)$ with $H_{2n-q}(X,\Z)$.
\end{proposition}

\begin{proof}
 Recall the definition that
$$\Phi:L_pH_{k+2p}(X)_{hom}\rightarrow \bigg\{\bigoplus_{{r> k+1,r+s=k+1}}
H^{p+r,p+s}(X)\bigg\}^*\bigg/H_{2p+k+1}(X,\Z)$$ is given by
$\Phi([f])=\Phi_f$, where $\Phi_f(\omega)=\int_{\tilde{c}}\omega
\quad ({\rm mod} \quad H_{2p+k}(X,\Z))$ with
$\partial(\tilde{c})=c_f$. The Lemma 12.10 in
\cite{Harvey-Lawson-Zweck} implies that the two constructions
coincide.
\end{proof}

\begin{remark}\label{sec:Rem3.4}
{ It is easy to see from the definition of $\Psi_*$  that the image of $\Phi_{p,{k+2p}}$ is in
$\ker(\Psi_*)$. Hence the natural map
$\Phi_{p,{k+2p}}:L_pH_{k+2p}(X)\rightarrow H_{k+2p}(X)\}$ factors through $\hat{a}$
and the map $\delta_2$ in the first $3\times3$ grid given in
\cite{Harvey-Lawson2}, p.26.}
\end{remark}

\begin{remark}
{ Gillet and Soul$\acute{e}$ \cite{Gillet-Soule} first showed
that the Griffiths' higher Abel-Jacobi map coincides with the
Riemannian Abel-Jacobi. Harris \cite{Harris} also discussed some
related topics.}
\end{remark}

In summary, we have the following

\begin{theorem}
{\label{sec:T7Th3.1} Let $X$ be a smooth complex projective variety
of dimension $n$. We have a well-defined map
$$\hat{a}: L_pH_{k+2p}(X)\rightarrow
H^{2(n-p)-k}_{\D}(X,\Z(n-p-k-1)),$$ given by
$$\hat{a}([f])=\widehat{a_f}
$$
which coincides with the generalized Abel-Jacobi map defined in
\cite{aj} when $\hat{a}$ is restricted on $L_pH_{k+2p}(X)_{hom}$; and the
projection of the image of $\hat{a}$ under $\delta_2$ is the natural
map $\Phi_{p,k}$. }
\end{theorem}

\begin{proof} The first statement follows from Lemma \ref{sec:T7L3.1} and \ref{sec:T7L3.2} that
$\hat{a}$ is well-defined. The second one follows from Proposition \ref{sec:T7Prop3.1} and the the third
one follows from \ref{sec:Rem3.4}.
\end{proof}

\begin{remark}
The map $\hat{a}$ can be nontrivial even if restricted on $L_pH_{k+2p}(X)_{hom}$.
This has been shown by Griffiths \cite{Griffiths} for $k=0$ and by the author
\cite{aj} for $k>0$.
\end{remark}

In this section below,  we will  show that the homomorphism $\hat{a}$ is well-behaviored with projective morphisms between
smooth projective varieties.

Recall that Friedlander and Lawson have defined  morphic cohomology groups  $L^pH^k(X)$ for all $k\leq 2p$.
Moreover,  they  have defined a duality between morphic
cohomology groups $L^pH^k(X)$ with Lawson homology groups $L_{n-p}H_{2n-k}(X)$
for a projective variety $X$ (cf. \cite{Friedlander-Lawson}, \cite{Friedlander-Lawson2}).

\begin{theorem}[\cite{Friedlander-Lawson2}]\label{duality}
If $X$ is smooth projective of $\dim X=n$, then the duality
$$\mathcal{D}: L^pH^k(X)\to L_{n-p}H_{2n-k}(X)$$
is an isomorphism for all $k\leq 2p$.
\end{theorem}

\begin{proposition}\label{prop3.10}
Let $\varphi:X\to Y$ be a projective morphism between smooth projective varieties $X$ and $Y$. Then we have the following
commutative diagram
$$
\xymatrix{&L^{p}H^{2p-k}(X)\ar[r]^-{\bar{a}_X}&H^{2p-k}\sb{\D}(X,\Z(p-k-1))&\\
&L^{p}H^{2p-k}(Y)\ar[u]^{\varphi^*}\ar[r]^-{\bar{a}_Y}&H^{2p-k}\sb{\D}(Y,\Z(p-k-1)).\ar[u]^{\varphi^*}&},
$$
where $\bar{a}_X= \hat{a}_X\circ\mathcal{D}$ and similarly for $\bar{a}_Y$.
\end{proposition}

\begin{proof} Let $n=\dim(Y)$. Given $\alpha\in L^{p}H^{2p-k}(Y)$, we set  $$\beta=\mathcal{D}(\alpha)\in L_{n-p}H_{2(n-p)+k}(Y)$$ and then choose a PL map $f:\rS^k\to \cZ_{n-p}(Y)$ such that $[f]=\beta$, 
where $[f]$ denotes the homotopy class of $f$.
Then we have $\bar{a}_Y(\alpha)=\hat{a}_Y([f])$.

 Let $c_f$ be the current constructed for $f$. On one hand, we have
\begin{equation}\label{eq1}
\varphi^*(\bar{a}_Y(\alpha))=\varphi^*(\hat{a}_Y\circ \mathcal{D}(\alpha))=\varphi^*(\bar{a}_Y([f]))=\varphi^*(-\Psi d^*G(c_f)).
\end{equation}

On the other hand,
\begin{equation}\label{eq2}
\begin{array}{llll}
\bar{a}_X(\varphi^*(\alpha))&=&\hat{a}_X(\mathcal{D}\varphi^*(\alpha))&\\
&=&\hat{a}_X(\varphi^*\mathcal{D}(\alpha))&\\
&=&\hat{a}_X(\varphi^*([f]))&\\
&=&\hat{a}_X([\tilde{\varphi}^*\circ f]),&\hbox{(cf. \cite{Peters}, p.218)}\\
&=&-\Psi d^*G(c_{\tilde{\varphi}^*\circ f}),&
\end{array}
\end{equation}
where $\tilde{\varphi}^*$ is Gysin homomorphism associated to the projective morphism $\varphi:X\to Y$ (cf. \cite{Friedlander-Gabber} for $\varphi$ a regular embedding, \cite{Peters} for any projective morphism).

 Note that $\varphi^*$ commutes with $\Psi$, $d^*$ and the Green operator $G$. by comparing Equation (\ref{eq1}) and (\ref{eq2}),  it is enough to show that
$$
\varphi^*(c_f)=c_{\tilde{\varphi}^*\circ f}
$$
which follows from the construction of $c_f$ and $c_{\tilde{\varphi}^*\circ f}$ (cf. \cite{aj}, Section 3). This completes the proof of the proposition.
\end{proof}

%%%%%%%%%%%%%%%%%444444444444444444444444444444444444444444444444444
\section {Relationships to other Abel-Jacobi type constructions}
In this section we discuss relations between our construction to other Abel-Jacobi type constructions,
e.g., the morphic Abel-Jacobi map constructed by M. Walker \cite{Walker}
and the Abel-Jacobi map for higher chow groups by M. Kerr,  J. Lewis and  S. M\"{u}ller-Stach  \cite{KLM}.

From  the construction of the homomorphism $\hat{a}$ in the last section,
we see that it essentially divided into two main steps.  First, for an element $\alpha \in L_pH_{k+2p}(X)$, we can find a PL
map $f:\rS^k\to \cZ_p(X)$ such that its homotopy class $[f]=\alpha$,
and  construct an integral cycle $c_f$ on $X$.
Then we associate this integral cycle a Hodge spark $\widehat{a_f}$ and show this spark is
in the suitable target space, i.e, the Deligne cohomology $H^{2(n-p)-k}_{\D}(X,\Z(n-p-k-1))$.

For the second step a Hodge decomposition is needed for currents on $X$.
Compact Riemannian manifolds are such examples.

The construction  in the first step depends on what we are interested in. In section \ref{sec3},
we deal with those integral currents from elements of Lawson homology. It can be applied to other cases.

Applying Example \ref{eg2.2} to $m=2(n-p)-k-1,q=n-p-k$, we get

\begin{proposition}\label{prop4.1}
Set $$\Omega^k(\cZ_p(X))_{htp\sim 0}:=\{f\in \Omega^k(\cZ_p(X))| \hbox{$f$ is PL and homotopically trivial} \}.$$
Set $b_f:= -\Psi_0 (d^*G(c_f))$, where  $\Psi_0$ denotes the projection corresponding to  $q=n-p-k$ in
Example \ref{eg2.2}. Then $(b_f,c_f)$ is the  Hodge spark of the map $f:{\rS}^k\rightarrow {\cZ}_p(X)$
and denote by  $\widehat{b_f}\in \widehat{\H}^{2(n-p)-k-1}(X,n-p-k)$ its associated
differential character.

Then we have a natural homomorphism
$$\hat{b}:\Omega^k(\cZ_p(X))_{htp\sim 0}\rightarrow
H^{2(n-p)-k-1}(X,n-p-k)/H^{2(n-p)-k-1}_{\Z}(X,n-p-k),$$ given by
$$\hat{b}([f])=\widehat{b_f}.
$$
\end{proposition}

\begin{proof}
What we need to show is $\delta_1(\widehat{b_f})=0\in \widehat{\H}^{2(n-p)-k-1}(X,n-p-k)$ since
$\ker(\delta_1)=H^{2(n-p)-k}_{\D}(X,\Z(n-p-k))$.
 By definition of $\delta_1$, we have $\delta_1(\widehat{b_f})=\widehat{\Psi_0(H(c_f))}$, where
$H(c_f)$ is the harmonic part of $c_f$. From the construction of $c_f$, we have
$$c_f\in \bigoplus_{r+s=k,|r-s|\leq k} {\D}'_{p+r,p+s}(X).$$

Note that $H(c_f)$ and $c_f$ are of the same $(*,*)$-type.
By the type reason, the projection of $H(c_f)$ under $\Psi_0$ on
$$\bigoplus_{r+s=k,r\geq k+1} {\D}'_{p+r,p+s}(X)
$$ is zero. This is exactly the image of $\widehat{b_f}$ under
$\delta_1$.

So we have a  homomorphism
$$\hat{b}:\Omega^k(\cZ_p(X))_{htp\sim 0}\rightarrow
H^{2(n-p)-k}_{\D}(X,\Z(n-p-k)),$$ given by
$$\hat{b}([f])=\widehat{b_f}.
$$

To compete the proof of the proposition, we need to show that $\delta_2(\widehat{b_f})=0$ since
$$\ker \delta_1\cap \ker \delta_2=H^{2(n-p)-k-1}(X,n-p-k)/H^{2(n-p)-k-1}_{\Z}(X,n-p-k).$$

From the definition of $\delta_2$, we have $\delta_2(\widehat{b_f})=[\Psi_0(c_f)]$, where $[c]$ denotes
the homological class of the integral cycle $c$. From the proof of Lemma \ref{sec:T7L3.2}, $c_f$ is the boundary of
an integral current $c_F$. Since the projection $\Psi_0$ commutes with the boundary map, we have $\Psi_0(c_f)=\partial (\Psi_0(c_F))$
and hence $[\Psi_0(c_f)]=0$. That is to say, $\delta_2(\widehat{b_f})=0$.
\end{proof}

Sometimes we denote by $J^k(X,p):=H^k(X,p)/H^k_{\Z}(X,p)$.  In propostion \ref{prop4.1},
those $c_f$  is called linearly equivalent to zero  if if $\widehat{b_f}=0$. Set
$\Omega^k(\cZ_p(X))_{0}:=\{f\in \Omega^k(\cZ_p(X))_{htp\sim 0}| \widehat{b_f}=0\}$. Therefore, we get an injection (also called $\hat{b}$)
on the quotient $$\overline{\Omega^k(\cZ_p(X))_{htp\sim 0}}:=\Omega^k(\cZ_p(X))_{htp\sim 0}/\Omega^k(\cZ_p(X))_{0}.$$
That is

$$
\hat{b}:\overline{\Omega^k(\cZ_p(X))_{htp\sim 0}}\longrightarrow J^{2(n-p)-k-1}(X,n-p-k).
$$

\begin{remark}
When $k=0$, the construction here concides with the Griffiths' Abel-Jacobi map on the space of algebraic cycles $p$-cycles which are algebraic equivalent to zero.
 (cf. \cite{Harvey-Lawson-Zweck}, Example 12.16).
\end{remark}

\begin{remark}
 The construction here for $k=0$ is different from the morphic Abel-Jacobi map constructed by Walker \cite{Walker}. On the one hand, we require $X$ is a \emph{smooth} complex \emph{projective} variety but there is no such requirement for these conditions in the definition.
On the other hand, at least in the case that $L_pH_{2p+1}(X)_{hom}\neq 0$, the morphic Jacobian $J^{mor}_p(X)$ defined in \cite{Walker} is different form $J^{2(n-p)-1}(X,n-p)$ defined above.
\end{remark}

For a smooth complex projective varieties $X$, S. Bloch \cite{Bloch}  introduced higher Chow groups  and constructed  regulator maps
$$
c\sb{p,l}:{\rm CH}^{p}(X,l)\rightarrow H^{2p-l}\sb{\D}(X,\Z(p))
$$
from the higher Chow groups to Deligne cohomology.  An explicit description of the restriction of $c_{p,l}$ on the homologically trivial part is given  by integration currents in \cite{KLM}. Here we descript a map from ${\rm CH}^{p}(X,l)$ to a different Deligne cohomology by using the system method in \cite{Harvey-Lawson-Zweck}.

Recall that (cf. \cite{Bloch}) for each $m\geq 0$, let
$$
\Delta[d]:=\{t\in \C^{d+1}|\sum_{i=0}^{m} t_i=1\}\cong \C^d.
$$
and let $z^l(X,d)$ denote the free abelian group generated by irreducible subvarieties of codimension-$l$ on $X\times \Delta[d]$ which
meets $X\times F$ in proper dimension for each face $F$ of $\Delta[d]$. Using intersection and pull-back of algebraic cycles,  we can define face and degeneracy relations and obtain a simplical abelian group structure for . Let $|z^l(X,*)|$ be the geometric realization of $z^l(X,*)$.  Then the higher Chow group is defined as
$$
\Ch^l(X,k):=\pi_k(|z^l(X,*)|).
$$

For $\alpha\in \Ch^l(X,k)$, let $f:\rS^k\to |z^l(X,*)|$ be a PL representative of $\alpha$. Since $\rS^k$ is compact, there is
an integer $d\geq 0$ such that $f(\rS^k)\subset \Delta[d]\times z^l(X,d)$. Note that for each $s\in \rS^k$, $f(s)$ is a codimension-$l$ algebraic cycles in $X\times \Delta[d]$ with certain additional restriction. As in section \ref{sec3},  we can construct an integral current $c_f$ on $X\times\Delta[d]$. Set $C_f:=pr_{X*}(c_f)$, where $pr_X$ is the projection on $X$. Note that $C_f$ is an integral current of suitable dimension since elements in $z^q(X,d)$ are required to intersect faces  properly.

Again  we consider Example \ref{eg2.2}  for $m=2p-k-1,q=p-k-1$. Set $A_f:=-\Psi (d^*G(C_f))$ and
Let $\widehat{A_f}\in \widehat{\H}^{2(n-p)-k-1}(X,n-p-k-1)$ be the
differential character corresponding to the Hodge spark $(A_f,C_f)$.

\begin{proposition}\label{prop4.4} Let $X$ be a smooth projective variety of dimension $n$.
 There is a homomorphism
$$\hat{A}\sb{p,k}:{\rm CH}^{p}(X,k)\rightarrow H^{2p-k}\sb{\D}(X,\Z(p-k-1)) $$
defined by
$$
\hat{A}\sb{p,k}([f])=\widehat{A_f}.
$$
\end{proposition}
\begin{proof}
Word for word to Lemma \ref{sec:T7L3.1}, we show that $\delta_1(\widehat{A_f})=0= \delta_2(\widehat{A_f})$. Similar to Lemma \ref{sec:T7L3.2}, $\widehat{A_f}$ is well-defined on ${\rm CH}^{p}(X,k)$.
\end{proof}

The combination of Theorem \ref{sec:T7L3.1} and Proposition \ref{prop4.4} gives us the following result.

\begin{corollary}\label{cor4.5}
 The range of  $\hat{A}\sb{p,k}$ constructed here  is different from the regulator map constructed by Bloch. From the construction,
we have a commutative diagram:
$$
\xymatrix{&\Ch^{p}(X,k)\ar[r]^-{FG}\ar[dr]^{\hat{A}\sb{p,k}}&L_{n-p}H_{2(n-p)+k}(X)\ar[d]^{\hat{a}}&\\
&&H^{2p-k}\sb{\D}(X,\Z(p-k-1))&},
$$
 where $FG$ is the Friedlander-Gabber homomorphism defined in \cite{Friedlander-Gabber}, Section 4.
\end{corollary}

\begin{center}{\bf
Acknowledgement}\end {center} I would like to express my gratitude to
my former advisor, Blaine Lawson,  for  all his help. I also  thank the referee
for detailed suggestions and corrections on the first version of the paper.

\end{document}